OpenStudy (anonymous):

i medal and fan. im asking for an example not a full answer please. i just put the full question is all. you have$60.00 in your wallet and want tp buy some new CDs. of the CDs are $11.00 each, what number of CDs, X, can you buy? write and solve an inequality. then explain your answer

OpenStudy (anonymous):

Example inequality: \[x \times z \le y\]Meaning that the product of x and z must be less than y.

OpenStudy (anonymous):

@fantasiablueskys see above

OpenStudy (anonymous):

thank you for the example @mrdoldum

OpenStudy (anonymous):

@fantasiablueskys Hope it helps.

OpenStudy (anonymous):

so the equation would be x*$11.00<_ $60.00

OpenStudy (anonymous):

ok so the equation would be X*$11.00<_ $60.00 then you would divide $60.00 by $11.00 and get 5.45454545...

OpenStudy (anonymous):

@fantasiablueskys Yes. \[X \times \$11.00\le\$60.00\]